A small mass slides down an inclined plane of inclination ‘s’ with horizontal.The coefficient of friction ‘D’ =D0x where x is the distance through which the mass slides down and D0 is a positive constant.find the distance covered by the mass before it comes to rest.

Grade:11

1 Answers

Vikas TU
7 years ago
Net Force = mgsin(s) – Dmgcos(s)
Net acceleration = gsin(s) – D0xgcos(s)
a = vdv/dx =  gsin(s) – D0xgcos(s)
      vdv =  gsin(s)dx – D0xgcos(s)dx
Integerating botth sides and with limit of v 0to 0.
v^2/2 = xgsin(s) – D0x^2gcos(s)/2
 xgsin(s) – D0x^2gcos(s)/2 = 0
Hence, x = (2/Do)(sins/cos(s)) =  > (2tan(s))/Do m.